# SAVeD: Learning to Denoise Low-SNR Video for Improved Downstream Performance

### v1

@@ Abstract
Foundation models excel at vision tasks in natural images but fail in low signal-to-noise ratio (SNR) videos, such as underwater sonar, ultrasound, and microscopy. We introduce Spatiotemporal Augmentations and denoising in Video for Downstream Tasks (SAVeD), a self-supervised method that denoises low-SNR sensor videos and is trained using only the raw noisy data. By leveraging differences in foreground and background motion, SAVeD enhances object visibility using an encoder-decoder with a temporal bottleneck. Our approach improves classification, detection, tracking, and counting, outperforming state-of-the-art video denoising methods with lower resource requirements. Project page: https://suzanne-stathatos.github.io/SAVeD/. Code page: https://github.com/suzanne-stathatos/SAVeD.
@@ 1 Introduction
Motion may be the only way to identify objects in video with low signal-to-noise-ratio (SNR), camouflage, or complex textures that may hinder frame-by-frame object detection. The human visual system is excellent at capturing observable motion [26], and this capability has not yet been reproduced by modern generative models. Learning to exploit motion cues will improve models’ ability to detect and track objects of interest in noisy video.
Obtaining sufficient annotations to train supervised models in video can be prohibitively expensive, especially for scientific [30] or medical [64, 15] applications, and delays the deployment of models when tackling novel signal statistics. Once trained, supervised detectors may not generalize well and may need additional annotations to be adapted to other locations with different background and foreground distributions[49], which may require additional supervision [14]. On the other hand, models trained using self-supervision can be more robust [60, 47]. This work aims to enhance motion signals in unlabeled low-SNR data, such as ultrasound and sonar videos, to improve downstream supervised classification, detection, and tracking.
Figure 1: SpatioTemporal Denoising improves classification, detection, tracking, and counting in video. We denoise sonar and ultrasound videos of fish in a river, lung scans, breast lesion scans, and cell microscopy to improve downstream classification, detection, tracking, and counting tasks. We propose a self-supervised method to enhance the foreground signal of video frames without manual annotations. Our method works on videos with: non-stationary backgrounds, low signal-to-noise-ratios, and a variable number of objects in a video.
Unsupervised and self-supervised methods are increasingly used for object localization and action recognition [38, 62, 16, 68, 58, 63]. However, existing methods do not address low-SNR videos. Furthermore, while some methods [21, 24] handle changes in camera view-angle, they do not handle cases where the camera is stationary and the background is not.
We address these challenges with SAVeD, a self-supervised learning method to denoise video. Additionally, we exploit object motion to boost the SNR across frames. Inspired by work on self-supervised reconstruction [38, 62], broad vs. narrow self-supervised video understanding [53], and anomaly detection [77, 45], we use an encoder to encode appearance frames, an hourglass network to combine temporal features, and a decoder network to reconstruct the denoised frame.
@@
  • We propose SAVeD, a novel denoising approach to clarify low-SNR video with variable numbers of agents; details are in Sec. 3.
  • We propose a rich benchmark for low-SNR video denoising consisting of a diverse collection of low-SNR video domains (sonar video of fish, ultrasound video of lungs and breast, microscopy video of tissue) and a diverse collection of downstream visual tasks (classification, detection, tracking, and counting), in Sec. 4.
  • We explore the value of different algorithmic choices in low-SNR video processing and test SAVeD, together with a number of variants (in Sec. 5 & A, and Supplementary materials), on our benchmark.
Figure 2: SAVeD, our approach for self-supervised denoising using spatiotemporal difference and identity reconstruction. $I_{t}$, $I_{t-T}$, and $I_{t-2T}$ are video frames at times t (current frame), t-T, and t-2T. These frames are input to an appearance encoder $\Phi$. The resulting feature representations are input to a spatiotemporal bottleneck $\Theta$ that compresses the 3 appearance features into a single spatiotemporal feature representation. Our model then predicts the reconstruction target, defined in Eq. 2 in Sec. 3.2, using the reconstruction decoder $\Psi$. The architecture is discussed in more detail in Sec. 3.3.
@@ 2 Related Work
Sonar and Ultrasound Sonar and ultrasound pose a distinct and interesting set of challenges to the computer vision community. Particularly noteworthy elements include: all sonar has pink noise in it, all non-cavity objects of interest have higher pixel intensity values compared to the background, and objects of interest may not be uniquely identifiable from their appearance features. Weld et al. [72] attempt to standardize ultrasound data through geometric analysis and augmentation, given ultrasound-data’s sensor variability. Unlike most computer vision datasets, which focus on light-based data, ultrasound, sonar, lidar, and radar do not rely on capturing light intensity. Instead, they rely on the principle of emitting waves, which bounce off objects, and return to the sensor as echoes. The “camera”, then, measures the distance to those objects by calculating the time it took for the echo to return. Sonar and ultrasound use sound waves, while lidar uses laser light pulses; sonar and ultrasound are primarily used in liquids while lidar is more often used for land and air-based applications [25, 9, 10]. We focus on one sonar dataset and two ultrasound datasets described in more detail in Sec. 4.1.
Classical image denoising. A large number of spatial filters from image processing techniques have been applied to image denoising [73, 67, 51, 4, 75, 65, 23] – they can be broken down to two types: linear and non-linear filters. Linear spatial filters. Mean and Gaussian filtering [27] reduces Gaussian noise, however, they can over-smooth noisy images [1]. Weiner filtering [4] aims to overcome this drawback, but it can overly blur sharp edges. Non-linear filters. With non-linear filters, noise can be reduced without first identifying the noisy pixels. Median-filtering [27] replaces each pixel with the median value of its neighboring pixels. Median and bilateral filtering [67] preserves edges while smoothing images to reduce the noise, though bilateral filtering is inefficient [23].
@@
Self-supervised and unsupervised image denoising. Several approaches use variants of blind-spot networks or pixel-wise masking to denoise imagery. Noise2Self [3] and Noise2Void (N2V) [40] train on noisy images without requiring clean targets or paired noisy data. N2V trains a blind-spot network to predict masked pixels’ intensity values based on neighboring pixels. Others [42, 3, 33] refrain from masking the pixels via a structural blind-spot network composed of half-plane receptive-field U-Nets [55]. Jang et al. [33] use a conditional blind-spot network and a loss that regularizes the denoised images without masking input pixels to train their network. Neighbor2Neighbor [31] proposes a self-supervised loss between two sub-sampled images. In general, noise in real-world imagery, including acoustic imagery, has unknown or non-stationary statistics that are spatially correlated, violating assumptions of pixel-wise independence.
CNN-based video denoising. Some video denoising methods leverage videos’ spatio-temporal structure by using optical flow for motion compensation [74, 66]. DVDnet [66] uses calculated flow-estimates to manually warp frames, align their contents, and process them collectively with a CNN. UDVD [59] uses a patch-wise noise-to-noise training strategy to predict clean frames by estimating masked pixels from adjacent neighborhoods of noisy frames.
@@
Denoising autoencoders (DAEs) were originally introduced to learn more robust representations. During training, DAEs intentionally add noise to their input data and learn to reconstruct the original uncorrupted signal. mDAE [20], a method for missing data imputation (replacing missing or unavailable data), improves performance on a handful of datasets. Zaki et al. [78] leverages a DAE framework as a preprocesing step to improve the quality of SERS spectra for biomarker quantification and discovery. They generated noisy data by duplicating background measurements at random locations. DAEs have also increasingly been applied to video tasks. CompDAE [50] explicitly models noise from snapshot compressive imaging measurements in low-light conditions to improve edge detection and depth estimation. TADA [12] uses an adversarial denoising autoencoder to remove EMG noise from EEG time series data. Our work similarly extends the application of DAEs to spatio-temporal sonar and ultrasound video denoising; we uniquely combine temporal frames to enhance signal quality while simultaneously addressing the increased noise introduced by this process.
Detection by tracking (DbT) methods locate objects in the first frame of a video and then track them to predict future locations [52, 2, 76, 48]. Point-tracking is similar, in that pixels are first initialized and then tracking models are trained to follow the pixels in a video [79, 17, 18, 70, 19, 39, 36, 35]. If the tracking is robust, detection is solved, even in frames where a detector alone fails. Tracking, therefore, helps fill the gap where detectors struggle. DbT is used in video object segmentation [52] to propagate masks over time, in long-term object tracking [34] which re-detects objects when needed, and in Visual Simultaneous Localization and Mapping, which relies on tracking features across frames to infer their presence or position of objects or the camera [11]. In an underwater setting, point-DbT for object discovery or detection may struggle due to the turbidity of water – discovering erraneous points since the background has motion. Inspired by DbT approaches, our method adds temporal information directly to the spatial dimension to boost the foreground signal.
Tracking by detection (TbD) [28, 13] approaches divide the tracking problem into two steps: first, an object detector predicts objects and their locations in every frame; then, a tracker associates detections from one frame to another, creating trajectories and uniquely identifying objects. While there has been notable progress in TbD recently [13], there are also particular drawbacks. TbD relies heavily on detection performance – false alarms, detection gaps, and missed object groupings may lead to incomplete tracks. In addition, some recent progress in TbD can be attributed to matching targets with spatially re-identifiable features. However, this is unrealiable in sonar. Kay et al. [37]’s downstream tracking is a tracking-by-detection approach; in order to improve it, then, we design our method to maximize downstream detection performance by lowering false positive and false negative rates concurrently.
@@ 3 Method
The goal of SAVeD (Fig. 2) is to remove noise while capturing and focusing on motion of objects of interest from video with a non-stationary, fluid background. Inspired by previous methods [62, 32, 57], we use an encoder-decoder setup. We propose a novel reconstruction target based on spatiotemporal differences in a neighborhood of frames. We rely on a background with a spatiotemporal distribution that is distinct from the foreground objects.
@@ 3.1 Self-supervised denoising
In low-SNR videos, signals are often distributed across multiple frames; as such, we want to condense information from multiple times into a single frame to exaggerate the signal. We do this through the reconstruction target. For simplicity, we choose to reconstruct the spatiotemporal combination of 3 frames, the current input frame $I_{t}$, the future frame $I_{t+T}$ and the previous frame $I_{t-T}$, from three input frames, $I_{t}$, $I_{t-T}$, and $I_{t-2T}$. We explore a vanilla autoencoder, UNets, and 3D convolutions (in Tab. 2), but ultimately find an encoder, bottleneck, and decoder framework works optimally.
We use an encoder-decoder architecture, seen in Fig. 2, with a spatial encoder, $\Phi$, a temporal hourglass network, $\Theta$, and a reconstruction decoder, $\Psi$. During training, spatial encoders, $\Phi$, takes $I_{t}$, $I_{t-T}$, and $I_{t-2T}$ as input to generate spatial feature embeddings, which are then used by the hourglass network, $\Theta$, to generate a spatiotemporal feature embedding; this embedding passes through $\Psi$ to reconstruct the learning objective $\hat{S}_{t,T}$.
@@ 3.2 Reconstruction target and loss
Target. We use the positive frame difference with the current frame (PFDwTN), which incorporates spatiotemporal information as our main reconstruction target.
This combines the current frame with the positive motion from the previous and next frames. Positive motion of the next frame is defined as $\max(0,I_{t}-I_{t+T})$, while positive motion from the previous frame is defined as $\max(0,I_{t}-I_{t-T})$. Note that the previous frame $I_{t-T}$ goes into the network, whereas the future frame $I_{t+T}$ does not. It is seen only when calculating the ground-truth target.
To handle frames where the background movement does not differ significantly from the foreground objects’ motion (i.e., stationary objects), we include the original frame, $I_{t}$, in the reconstruction target. These signals boost the spatial signal by exploiting the motion signature. The overall target is:
$S_{t,T}=\max(0,I_{t}-I_{t-T})+I_{t}+\max(0,I_{t}-I_{t+T})$ (2)
Other motion-augmenting targets that we tested are defined and visualized in Sec. A.1 and Fig. 9 in the Supplemental materials.
@@ 3.3 Noise Removal Network
Appearance Encoder $\Phi$. We implement a 6-layer CNN. Each layer consists of a convolutional block (Conv2D + ReLU) followed by max pooling, progressively increasing the number of feature channels while reducing the spatial dimension, $\mathbb{R}^{(H,W,1)}\rightarrow\mathbb{R}^{(\frac{H}{32},\frac{W}{32},512)}$. We also save skip connections, which are sequential max pools followed by 1x1 convolutions, to be used by the hourglass network and decoder. This design mimics the encoding portion of a UNet, with a fraction of the parameters and FLOPS, to let the network capture multi-scale features efficiently.
@@
More details can be seen on each of these in the Supplemental materials Tab. 6.
We recognize that combining noisy frames adds to the noise of the overall signal rather than removing it. Work [5] has shown that denoising methods capture clean data’s underlying structure. Denoising autoencoders purposefully corrupt input data by adding noise or masking some of the input values [68, 29]. We rely on the autoencoder to remove noise implicitly by focusing on the largest reconstruction areas to minimize loss. This assumes that the objects of interest are larger than the noise signature.
@@ 3.4 Denoising Metric
Typically [40, 42, 33, 31, 59], denoising networks use Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) [71] as evaluation metrics. PSNR is the ratio of maximum signal power to noise power, and SSIM measures perceived image quality. Both metrics rely on having clean imagery to compare with. Our denoising approach is unsupervised and we do not have clean imagery. As a result, we design a pseudo-PSNR metric to indicate denoised performance for object detection, and we rely on downstream performance for other tasks.
3.4.1 Pseudo-PSNR for Downstream Detection
Recall that we assume that our downstream models are supervised. Therefore, we can assume we have bounding boxes or segmentation masks for detection tasks. For simplicity, we call detection annotations “boxes”, though the same approach works for segmentation masks. For each object’s box $b$ in each frame $I$ of each video, we take the density of pixel intensity values: $d_{b}=I[b]$, where $b$ are the indices associated with the box. Then, we take the density of pixel intensity values from a different frame $\tilde{I}$ of the same video at the same box location $b$ where we know there is no object, $\tilde{d}_{b}$. These densities give us pseudo-distributions between objects (signals) and background (noise). We postulate that if the distributions are separable i.e., the distribution of object pixels is distinct from the distribution of background pixels, then the denoising method works as intended. Therefore, we calculate the distance between object and non-object via the Kullback-Leibler (KL) Divergence [41]. KL divergence measures the distance between two probability distributions P and Q as follows:
[…]

### v2

@@ Abstract
Low signal-to-noise ratio (SNR) videos—such as those from underwater sonar, ultrasound, and microscopy—pose significant challenges for computer vision models, particularly when paired clean imagery for denoising is unavailable. We present Spatiotemporal Augmentations and denoising in Video for Downstream Tasks (SAVeD), a novel self-supervised method that denoises low-SNR sensor videos using only raw noisy data. By leveraging distinctions between foreground and background motion and exaggerating objects with stronger motion signal, SAVeD enhances foreground object visibility and reduces background and camera noise without requiring clean video. SAVeD has a set of architectural optimizations that lead to faster throughput, training, and inference than existing deep learning methods. We also introduce a new denoising metric, FBD, which indicates foreground-background divergence for detection datasets without requiring clean imagery. Our approach achieves state-of-the-art results for classification, detection, tracking, and counting tasks and it does so with fewer training resource requirements than existing deep-learning-based denoising methods. Project page here, Code: https://github.com/suzanne-stathatos/SAVeD.
@@ 1 Introduction
Motion may be the only way to identify objects in video with low signal-to-noise-ratio (SNR), camouflage, or complex textures that may hinder frame-by-frame object detection. The human visual system is excellent at capturing observable motion [15], a capability which has not yet been reproduced by modern computer vision models. Learning to exploit motion cues will improve models’ abilities to detect and track objects of interest in noisy video.
In several scientific [18] and medical applications [48, 10], clean (noise-free) imagery are not available to train image or video denoisers. One attractive element of self-supervised models is their ability to find the signal within the noisy imagery itself. Self-supervision can, too, be more robust and generalizable to various noise type [45, 36].
This work aims to enhance motion signals in low-signal-to-noise (SNR) data with non-stationary backgrounds, such as underwater, ultrasound, and sonar videos, to improve downstream supervised classification, detection, and tracking. We address sensor, background, and noise challenges with SAVeD, a self-supervised learning method to boost signal in noisy video. We exploit object motion to boost the SNR across frames. Inspired by work on self-supervised reconstruction [25, 47], self-supervised video understanding [39], and anomaly detection [59, 33], we use an encoder to encode appearance frames, a temporal bottleneck, and a decoder network to reconstruct the denoised frame.
Figure 1: SpatioTemporal Denoising improves classification, detection, tracking, and counting in video. We denoise sonar and ultrasound videos of fish in a river, lung scans, breast lesion scans, and cell microscopy to improve downstream classification, detection, tracking, and counting tasks. We propose a self-supervised method to enhance the foreground signal of video frames without manual annotations or clean imagery. Our method works on grayscale videos with: non-stationary backgrounds, low signal-to-noise-ratios, and a variable number of objects in a video.
@@
  • We propose SAVeD, a novel, state-of-the-art, resource-efficient self-supervised method to improve the signal in low-SNR video with variable numbers of agents and non-stationary background; details are in Sec. 3.
  • We introduce a novel denoising metric, $FBD$, that does not rely on clean imagery; it instead relies on background-foreground appearance distribution differences, in Sec. 3.
  • We propose a rich benchmark for low-SNR video denoising consisting of a diverse collection of low-SNR video domains (sonar video of fish, ultrasound video of lungs and breast tissue, and microscopy video) and a diverse collection of downstream visual tasks (classification, detection, tracking, and counting), in Sec. 4.
Figure 2: SAVeD, our approach for self-supervised denoising using spatiotemporal difference and identity reconstruction. $I_{t}$, $I_{t-T}$, and $I_{t-2T}$ are video frames at times t (current frame), t-T, and t-2T. These frames are input to an appearance encoder $\Phi$. The resulting feature representations are input to a spatiotemporal bottleneck $\Theta$ that compresses the 3 appearance features into a single spatiotemporal feature representation. Our model then predicts the reconstruction target, defined in Eq. 2 in Sec. 3.2, using the reconstruction decoder $\Psi$. The loss, defined in Eq. 3, is calculated and backpropagated through all networks. The architecture is discussed in more detail in Sec. $3.$
@@ 2 Related Work
Several works in the image and video denoising space require paired clean/noisy imagery for training, evaluation, or both, limiting their applicability in real-world denoising scenarios [60, 30, 23, 35]. We, notably, do not, and we focus this section on existing self-supervised and unsupervised approaches.
@@
Self-supervised and unsupervised image (frame-by-frame) denoising. Several approaches use variants of blind-spot networks or pixel-wise masking to denoise imagery, including frame-by-frame video. Noise2Self [2] and Noise2Void (N2V) [26] train on noisy images without requiring clean targets or paired data. N2V trains a blind-spot network to predict masked pixels’ intensities from neighboring pixels. Others [28, 2, 22] refrain from masking pixels via structural blind-spot networks with half-plane receptive-field U-Nets [41]. Jang et al. [22] use a conditional blind-spot network and a loss that regularizes denoised images without masking input pixels. Neighbor2Neighbor [19] proposes a self-supervised loss between two sub-sampled images. In general, noise in real-world imagery, including acoustic imagery, has unknown or non-stationary statistics that are spatially correlated, violating assumptions of pixel-wise independence.
Video Denoising with Flow Estimation Some video denoising methods leverage videos’ spatiotemporal structure by using optical flow for motion compensation [56, 49]. DVDnet [49] uses calculated flow-estimates to manually warp frames, align their contents, and process them collectively with a CNN. VDFlow [57] jointly learns video deblurring and optical flow. These approaches target video where one component is moving (e.g. an object or the camera). However, when objects’ motion affects the background’s motion, or when objects are small, we find that leveraging optical flow enlarges objects’ motion and makes it difficult to distinguish multiple closely-located objects. An example is in Fig. 16 in the Supp Mat.
Video Restoration and Denoising with Spatial and Temporal Consistency UDVD [44] uses a patch-wise noise-to-noise training strategy to predict clean frames by estimating masked pixels from adjacent neighborhoods of noisy frames. It relies on temporal redundancy across frames and does not rely on clean imagery, making it appealing for in-the-wild data. However, UDVD assumes gaussian noise, while underwaters sonar and ultrasound videos have pink and speckle noise [58, 20, 37]. UDVD’s patchwise nature may cause the model to overwhelmingly learn background when most of the video is spatially and temporally dominated by background. UMVD [1] extends UDVD to focus on microscopy data. It uses frame-level augmentations and a reconstruction loss that predicts each frame from its noisy temporal neighbors. It assumes that the signal is consistent across frames while the noise is not, and it assumes smooth object motion. However, like UDVD, it is also patch-wise.
AverNet [61], another self-supervised video restoration model, similarly relies on temporal consistency to fix time-varying unknown degradations. It uses two modules: 1.) prompt-guided alignment to line up video frames at the pixel level and 2.) prompt-conditioned enhancement, which restores each frame by adapting to the image’s degredation. The method was tested and performed well on videos with varying levels of gaussian, poisson, and speckle noise, gaussian and resizing blur, and jpeg and video compression.
LG-BPN [54], another patch-wise denoising method, uses two main ideas: local guidance and internal consistency. Local guidance preserves image structures like edges or textures that are consistent even with noise. Internal consistency assumes nearby patches in space and time share similar structure. However, LG-BPN struggles with long-range temporal dependency and highly-structured noise, such as sensor artifacts. RVRT [31] is a transformer-based model for denoising, deblurring, and super-resolution. It processes videos frame-by-frame, but models long-term dependencies (remembering what it learned from frames that are further apart) using spatiotemporal attention. It, however, requires very large memory even for inference.
@@
Denoising autoencoders (DAEs) were originally introduced to learn robust representations. During training, DAEs intentionally add noise to input data and learn to reconstruct the original uncorrupted signal. mDAE [11], a method for missing data imputation (replacing missing or unavailable data), improves performance on several datasets. DAEs have also been applied to video tasks. CompDAE [38] explicitly models noise from snapshot compressive imaging measurements in low-light conditions to improve edge detection and depth estimation. TADA [9] uses an adversarial denoising autoencoder to remove EMG noise from EEG time series data. Our work similarly extends the application of DAEs to spatio-temporal grayscale video denoising; we uniquely combine temporal frames to enhance signal quality while simultaneously addressing the increased noise introduced by this process.
Sonar and Ultrasound Sonar and ultrasound present challenges to the computer vision community. Key characteristics include pink noise, brighter pixel intensities for non-cavity objects compared to the background, and limited distinctiveness from appearance features. Weld et al. [55] address ultrasound sensor’s variability via geometric analysis and augmentation. Unlike many computer vision datasets based on signal from light intensity, ultrasound, sonar, lidar, and radar rely on wave echoes. The “camera” emits waves that reflect off objects and return to the sensor; distance is measured by the echo’s return time. Sonar and ultrasound use sound waves (mainly in liquids), while lidar uses laser pulses and is common in air and land environments [14, 7, 8]. We focus on one sonar dataset and two ultrasound datasets described in more detail in Sec. 4.1.
@@ 3 Method
The goal of SAVeD (Fig. 2) is to enhance objects’ signal by isolating and emphasizing their motion in video with a non-stationary, fluid background. Inspired by prior work [47, 21, 43], we use an encoder-decoder framework. Our contribution is twofold: (1) a reconstruction target based on spatiotemporal differences across neighboring frames, and (2) a denoising metric that does not require clean ground truth. The method assumes that background spatiotemporal statistics differ from those of the foreground.
@@ 3.1 Self-supervised denoising
In our benchmarks’ low-SNR videos, signals are distributed across time; as such, we want to condense information from multiple timesteps into a single frame to exaggerate the signal. We do this through the reconstruction target. For simplicity, we choose to reconstruct the spatiotemporal combination of 3 frames, expressed, in Eq. 2, from three input frames, $I_{t}$, $I_{t-T}$, and $I_{t-2T}$. We explore a vanilla autoencoder, UNets with and without skip connections and residual layers, and 3D convolutions (in Tab. 3), and find a simple encoder-bottleneck-decoder framework optimal. We also explore $N>3$ input frames, $T>1$, the target without the DAE network as input to downstream tasks, and more in our ablations in the Supplemental Tab. 4, 5, and Fig. 10.
We use an encoder-decoder architecture, seen in Fig. 2, with a spatial encoder, $\Phi$, a temporal hourglass network, $\Theta$, and a reconstruction decoder, $\Psi$. During training, spatial encoders, $\Phi$, take $I_{t}$, $I_{t-T}$, and $I_{t-2T}$ as input to generate spatial feature embeddings, which are then used by the hourglass network, $\Theta$, to generate a spatiotemporal feature embedding; this embedding passes to $\Psi$ to reconstruct the learning objective $\hat{S}_{t,T}$.
@@ 3.2 Reconstruction target and loss
Target. We use the directionally-positive frame difference with the current frame (PFDwTN) as our reconstruction target. This combines the current frame with the positive motion from the previous and next frames. Directionally-positive motion of the next frame is defined: $\max(0,I_{t+T}-I_{t})$; of the previous frame, it is defined: $\max(0,I_{t-T}-I_{t})$. In both cases, motion is relative to the current frame $I_{t}$. Note that the previous frame $I_{t-T}$ goes into the network, while the future frame $I_{t+T}$ does not. It is used only when calculating the ground-truth target for the loss.
To handle frames where the background movement does not differ significantly from the foreground objects’ motion (i.e., stationary objects), we include the original frame, $I_{t}$, in the reconstruction target. The overall target is:
$S_{t,T}=\max(0,I_{t-T}-I_{t})+I_{t}+\max(0,I_{t+T}-I_{t})$ (2)
Other motion-augmenting targets that we tested are defined/visualized in Sec. A.1 and Fig. 10 in the Supp. Mat.
@@ 3.3 Noise Removal Network
Appearance Encoder $\Phi$. We implement a 6-layer CNN. Each layer consists of a convolutional block (Conv2D + ReLU) followed by max pooling, progressively increasing the number of feature channels while reducing the spatial dimension, $\mathbb{R}^{(H,W,1)}\rightarrow\mathbb{R}^{(\frac{H}{32},\frac{W}{32},512)}$. We also save skip connections (sequential max pools and 1x1 convolutions) to be used by the hourglass network and decoder. This design uses a fraction of the parameters and FLOPs from an off-the-shelf UNet, to let the network capture multi-scale features efficiently. See Tab. 1 for efficiency comparisons.
@@
More details on each of these modules is in the Supplemental materials Tab. 6.
@@ 3.4 Denoising Metric
Typically [26, 28, 22, 19, 44], denoising networks use Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) [53] as evaluation metrics. PSNR is the ratio of maximum signal power to noise power, and SSIM measures perceived image quality. Both metrics rely on having clean imagery for comparison. Our denoising approach is unsupervised – we do not have clean imagery. As a result, we design a new metric, Foreground-to-Background Divergence ($FBD$), to evaluate denoised performance when we have detection annotations, and we rely on downstream performance as a proxy for denoised performance on datasets with different task annotations.
3.4.1 Foreground-to-Background Divergence (FBD) Metric for Unsupervised Denoising
Recall that we assume that our downstream models are supervised. Therefore, for detection tasks, we can assume we have bounding boxes or segmentation masks. For simplicity, we call detection annotations “boxes”, though the same approach works for segmentation masks.
FBDFBD measures how well a denoising method makes objects distinguishable from background by computing the KL-Divergence between a region containing an object (foreground) and the same region at a different time without it (background). An example is shown in Fig. 3. Unlike PSNR, which requires a clean reference, $FBD$ does not, which is critical in many real-world applications.
Figure 3: High-SNR visualization and calculations for FBDFBD and PSNR on Fly-vs-Fly [13] with synthetic noise. In order to calculate the PSNR of (B), a ‘clean’ image (A) is needed. (A) is not needed for FBDFBD, though (C) is.
For each object’s box $b$ in each frame $I$ of each video, we take the density of pixel intensity values: $d_{b}=I[b]$, where $b$ are the indices associated with the box. Then, we take the density of pixel intensity values from a different frame $\tilde{I}$ of the same video at the same box location $b$ where we know there is no object, $\tilde{d}_{b}$. If the distributions are separable i.e., the distribution of object pixels is distinct from the distribution of background pixels, then the denoising method works as intended. We measure the separability between object and non-object via the Kullback-Leibler (KL) Divergence [27]. KL divergence measures the distance between two probability distributions P and Q as follows:
$D_{KL}(P\|Q)=\int p(x)\log(\frac{p(x)}{q(x)})dx$ (4)